A sample of C2H4 with a pressure of 1.78 atm and a volume of 295 mL is allowed to react with excess H2 at 151 °C. H2 (g) + C2H4 (g) C2H6 (g)

Calculate the pressure of the C2H6 produced in the reaction if it is transferred to a 2.62-L flask and cooled to 34 °C.
Answer #1

Step 1: calculate number of mol of C2H4

Given:

P = 1.78 atm

V = 295.0 mL

= (295.0/1000) L

= 0.295 L

T = 151.0 oC

= (151.0+273) K

= 424 K

find number of moles using:

P * V = n*R*T

1.78 atm * 0.295 L = n * 0.08206 atm.L/mol.K * 424 K

n = 1.509*10^-2 mol

Step 2: Calculate number of mol C2H6

The balanced reaction is:

H2 (g) + C2H4 (g) —> C2H6 (g)

So,

mol of C2H6 produced = mol of C2H4

= 1.509*10^-2 mol

Step 3: calculate pressure of C2H6

Given:

V = 2.62 L

n = 1.509E-2 mol

T = 34.0 oC

= (34.0+273) K

= 307 K

use:

P * V = n*R*T

P * 2.62 L = 0.0151 mol* 0.08206 atm.L/mol.K * 307 K

P = 0.1451 atm

Answer: 0.145 atm